# Can Large Language Models Recall Reference Location Like Humans?

### v1

@@ Abstract
When completing knowledge-intensive tasks, humans sometimes need not just an answer but also a corresponding reference passage for auxiliary reading. Previous methods required obtaining pre-segmented article chunks through additional retrieval models. This paper explores leveraging the parameterized knowledge stored during the pre-training phase of large language models (LLMs) to independently recall reference passage from any starting position. We propose a two-stage framework that simulates the scenario of humans recalling easily forgotten references. Initially, the LLM is prompted to recall document title identifiers to obtain a coarse-grained document set. Then, based on the acquired coarse-grained document set, it recalls fine-grained passage. In the two-stage recall process, we use constrained decoding to ensure that content outside of the stored documents is not generated. To increase speed, we only recall a short prefix in the second stage, then locate its position to retrieve a complete passage. Experiments on KILT knowledge-sensitive tasks have verified that LLMs can independently recall reference passage location in various task forms, and the obtained reference significantly assist downstream tasks.
Can Large Language Models Recall Reference Location Like Humans?
@@ 1 Introduction
Although LLMs have impressive capabilities, generating complete reference passage may be too time-consuming. To address this issue, we explore whether LLMs can locate passage by recalling a prefix only. First, we prompt the LLM to generate a short prefix, then use the FM-index to locate the document containing this prefix within the document set. Finally, we use the Knuth-Morris-Pratt (KMP) algorithm to identify the starting position of the prefix in the document, determining a longer passage as the final reference from this position. We find that this method significantly speeds up the LLM’s recall of reference. And the location of such a short prefix still maintains a high quality of reference retrieval. Our framework requires no additional retrieval models and pre-segmentation, allowing the LLM to independently and naturally obtain natural reference from articles of any length, ultimately naming it LLMRefLoc.
We conduct comprehensive experiments on 6 knowledge-sensitive tasks from the KILT benchmark (Petroni et al., 2021). We find that in a zero-shot setting, merely prompting the open-source LLMs Llama (Touvron et al., 2023a) and Llama-2 (Touvron et al., 2023b) enables the effective recall of needed page documents and fine-grained reference passage under the LLMRefLoc framework, enhancing the performance of downstream tasks. In summary, our main contributions are as follows:
@@
Figure 2: Shows the LLMRefLoc framework. First, all Wikipedia titles are stored in a prefix tree, then the LLM is prompted to recall title identifiers under this prefix tree constraint. Subsequently, an FM-index is constructed from the top $k$ documents obtained, and the LLM recalls reference passage under the new constraint. The grey area represents the short prefix, used for the next step of locating and extracting the complete passage.
2 LLMRefLoc
In this section, we detail our two-stage framework, LLMRefLoc. In the first stage, we prompt the LLM to recall title identifiers, which serve as candidate documents for the next stage. In the second stage, the LLM is prompted to recall reference passage from the documents obtained in the first stage. To increase speed, we only recall a short prefix, then locate and extract the reference within the document. The structure is shown in Figure 2.
@@ 2.3 Short Prefix Recall and Localization
Subsequently, we find the document $d$ corresponding to $p_{s}$ within the document set $D_{k}$ obtained in the first stage. Since we have controlled the number of documents in the document set $D_{k}$ obtained in the first stage, in most cases, we can obtain a unique document $d$ containing $p_{s}$. When we cannot obtain a unique document, we simply default to selecting the first document. In the next step, we use the Knuth-Morris-Pratt (KMP) string matching algorithm to quickly determine the starting position $st$ of $p_{s}$ in $d$, then extract a complete reference $p_{final}=d[st:st+l_{p}]$ starting from $st$, with a length of $l_{p}$ tokens. The autoregressive score of the short prefix is also calculated using Equation 2 and the final two-stage score is obtained using Equation 3 to select the final reference. In Experiment 3.2.1, we find that recalling just the prefix for localization still achieves good results in obtaining reference.
@@ 3.1 Experimental Setup
${}^{\star}$ 54.74 45.68 25.46 $\mathbf{16.19}$ 56.61 26.62
LLMRefLoc(Llama-7b) 54.46 $\mathbf{57.03}$ 44.56 15.13 76.57 52.91
LLMRefLoc(Llama-13b) 54.42 55.53 46.30 12.94 77.55 34.51
LLMRefLoc(Llama-2-7b) 56.33 56.43 46.20 14.60 77.29 49.61
LLMRefLoc(Llama-2-13b) $\mathbf{57.77}$ 54.41 $\mathbf{48.70}$ 15.00 $\mathbf{83.69}$ $\mathbf{57.63}$
@@
${}^{\star}$ $\mathbf{47.94}$ 66.60 20.29 14.40 41.22 45.38
LLMRefLoc(Llama-7b) 36.87 58.48 25.55 15.99 54.85 59.40
LLMRefLoc(Llama-13b) 37.72 60.96 26.34 14.80 55.20 50.79
LLMRefLoc(Llama-2-7b) 38.07 62.88 27.55 $\mathbf{16.85}$ 56.23 57.79
LLMRefLoc(Llama-2-13b) 40.82 $\mathbf{68.20}$ $\mathbf{30.04}$ 15.06 $\mathbf{58.42}$ $\mathbf{63.43}$
@@
${}^{\star}$ $\mathbf{33.49}$ 72.68 23.13 $\mathbf{20.75}$ 75.27 14.17
LLMRefLoc(Llama-7b) 29.78 70.18 24.61 20.60 78.10 14.47
LLMRefLoc(Llama-13b) 29.68 71.60 25.48 20.24 78.53 14.33
LLMRefLoc(Llama-2-7b) 29.89 70.04 25.55 20.50 78.04 14.48
LLMRefLoc(Llama-2-13b) 31.69 $\mathbf{72.94}$ 26.13 20.61 $\mathbf{78.79}$ $\mathbf{14.77}$
@@ 3.2.1 Page-level Results
Coarse-grained page-level results, as shown in Table 1, demonstrate that the LLMRefLoc framework, when implemented with Llama-2-13b, achieves the best R-precision scores of 57.77, 48.70, 83.69, and 57.63 on the NQ, HotpotQA, FEVER, and WoW datasets, respectively. This significantly surpasses the performance of sparse retrieval BM25 and dense retrieval Contriever in a zero-shot scenario. It also shows strong competitive power against the fully trained DPR method, especially on the WoW and FEVER datasets, with improvements of 27.08 and 31.01 points, respectively. This result is consistent with the hypothesis that LLMs are powerful in recalling coarse-grained title identifiers, enabling the acquisition of high-quality relevant pages that assist in the subsequent fine-grained recall stage.
@@ 3.2.2 Passage-level Results
Fine-grained reference passage results, as shown in Table 2, reveal that the LLMRefLoc framework, when implemented with Llama-2-13b, also achieves the best scores of 68.20, 30.04, 58.42, and 63.43 on the TriviaQA, HotpotQA, FEVER, and WoW datasets, respectively. We note that the improvement of the framework in fine-grained reference passage compared to the DPR method is relatively reduced compared to the page-level results. This suggests potential for optimization in activating LLMs to recall more detailed and longer reference, presenting a greater challenge compared to recalling shorter title. Notably, DPR performs excellently on the NQ dataset, which may relate to its training data format. Interestingly, in the HotpotQA dataset, BM25 remains competitive, surpassing dense retrieval methods, possibly due to the longer questions in this dataset leading to more vocabulary overlap. LLMRefLoc shows significant progress on the FEVER and WoW datasets, demonstrating the potential and adaptability of LLMs in recalling high-quality reference passage across different task formats. Furthermore, the general enhancement in performance with the progression from Llama to Llama-2 and the increase in model size indicates a correlation between the recall ability and the underlying capabilities of LLMs.
@@ 3.2.3 Downstream Task Results
Downstream task results are presented in Table 3. LLMRefLoc, based on Llama-2-13b recalled passage, achieved the best scores of 72.94, 78.79, and 14.77 on the TriviaQA, FEVER, and WoW downstream tasks, respectively, validating the performance of LLM recall references in downstream tasks. On the open-domain question answering NQ dataset, although DPR performed excellently after full data training, LLMRefLoc also displayed highly competitive performance. On the other hand, in the TriviaQA and HotpotQA datasets, due to the length of the questions, BM25 achieved excellent performance by obtaining more vocabulary overlap, yet LLMRefLoc still achieved comparable or better performance in most cases. The unsupervised trained Contriever performed relatively poorly across all tasks, emphasizing the crucial role of supervised training in enhancing the performance of dense retrieval models.
@@ 3.4 Further Analysis
Memory Usage Analysis Dense retrieval methods such as Contriever and DPR require over 60GB of memory usage. In contrast, sparse retrieval methods use far less memory, only needing 17GB. The LLMRefLoc framework, utilizing FM-index and Trie indexing, requires only 8GB when pre-encoding and storing all documents with FM-index, and the Trie storing all title identifiers needs just 25MB, which is negligible. Compared to sparse and dense retrieval methods, the recall framework effectively saves memory.
@@ 5 Conclusion and Future Work
This paper introduces a framework named LLMRefLoc, which utilizes LLMs to independently recall reference passage and can be flexibly applied to various knowledge-sensitive tasks. Mimicking the human habit of searching for information, we first prompt the LLM to recall relevant document pages, and then from these pages, recall and locate reference passage. Furthermore, through beam search constrained by Trie and FM-index, we ensure that the content recalled by the LLM is a subset of existing texts. This framework can be flexibly used with various open-source LLMs, expanding the potential applications of LLMs. In the future, we consider enhancing the ability of LLMs to recall under constraints through instruction tuning, as well as applying this framework to more retrieval fields, exploring lightweight methods to inject new document knowledge into LLMs, and integrating multi-hop reasoning into reference recall.
@@
Although LLMRefLoc demonstrates the potential of LLMs to recall reference passage in knowledge-sensitive tasks like humans, its application still faces several limitations. Firstly, this framework struggles to surpass the performance of the current state-of-the-art retrieval models, especially those models that have been fine-tuned on specific tasks through supervision. In the future, there is a need to explore more effective ways of instruction tuning for recalling under constraints. At the same time, LLMRefLoc relies on document title identifiers for phased recall, meaning that its recall capability may be limited for documents lacking clear titles or identifiers.
Moreover, the framework finds it challenging to effectively recall documents that appear less frequently in the pre-training stage. This indicates that if a document appears infrequently in the training data of the LLM, or if the document content significantly differs from the training data, LLMRefLoc may encounter difficulties in recalling these documents. For the updating of documents and the injection of new knowledge, LLMRefLoc requires additional training to inject this new information into the model parameters. There is still a need to explore more efficient, lightweight methods for injecting new documents in the future.
@@ C.1 Prompts for the First Stage
LLMRefLoc(Llama-7b) 54.46 57.03 44.56 36.87 58.48 25.55
LLMRefLoc(Vicuna-1.3-7b) 48.47 47.99 40.79 35.28 56.41 23.75
LLMRefLoc(Llama-13b) 54.42 55.53 46.30 37.72 60.96 26.34
LLMRefLoc(Vicuna-1.3-13b) 52.73 46.61 43.41 36.55 67.29 26.63
LLMRefLoc(Llama-2-7b) 56.33 56.43 46.20 38.07 62.88 27.55
LLMRefLoc(Llama-2-chat-7b) 3.31 1.12 0.98 4.09 3.97 3.43
LLMRefLoc(Vicuna-1.5-7b) 50.76 51.73 41.23 34.16 55.98 24.43
LLMRefLoc(Llama-2-13b) 57.77 54.41 48.70 40.82 68.20 30.04
LLMRefLoc(Llama-2-chat-13b) 1.94 1.60 1.55 6.38 7.93 4.71
LLMRefLoc(Vicuna-1.5-13b) 52.24 56.34 45.90 37.22 63.24 27.14
@@ Appendix D Experimental Results for Different Values of Alpha
In Figure 4, we compare the experimental results of LLMRefLoc when implemented based on Llama-2-13b with different $\alpha$ values. When $\alpha=0.0$, it’s equivalent to not having a two-stage weighted method, relying only on the scores from the second stage’s fine-grained passage recall, resulting in the selection of suboptimal reference. With the increase of $\alpha$, the model sees improvements in both page-level and passage-level results, proving the importance of the first stage document scores for the final reference selection. However, when $\alpha$ reaches 0.95 and continues to increase, the final performance actually decreases to some extent, indicating the need to find a balance between the two for better results.
@@ Appendix E Experimental Results for Different Prefix Lengths
In Figure 5, we conduct experiments with LLMRefLoc recalling different numbers of prefix tokens based on Llama-2-13b. We observe that longer prefix lengths do not bring additional performance improvements; on the contrary, they lead to a decrease in performance. Existing LLMs still perform better when generating shorter passages under constraints; longer passages introduce additional noise, resulting in decreased performance. However, overly short prefixes might also lack sufficient information, leading to an inability to accurately select the desired passage as a reference.
@@ Appendix G Experimental Results under Few-Shot Prompts
In Figure 6, we compare the experimental results of LLMRefLoc when incorporating few-shot prompts in the second stage of fine-grained passage recall based on Llama-2-13b. Using more sample prompts only brings partial improvements on the HotpotQA dataset, and the degree of improvement tends to stabilize when the number of samples increases from 3 to 5. On the NQ and TriviaQA datasets, no further improvements are observed, and a slight decline is even noted. There is still a need to explore more effective prompting methods.

### v2

# Generative Retrieval with Large Language Models

@@ Abstract
When completing knowledge-intensive tasks, humans sometimes need not just an answer but also a corresponding reference passage for auxiliary reading. Previous methods required obtaining pre-segmented article chunks through additional retrieval models. This paper explores leveraging the parameterized knowledge stored during the pre-training phase of large language models (LLMs) to independently recall reference passage from any starting position. We propose a two-stage framework that simulates the scenario of humans recalling easily forgotten references. Initially, the LLM is prompted to recall document title identifiers to obtain a coarse-grained document set. Then, based on the acquired coarse-grained document set, it recalls fine-grained passage. In the two-stage recall process, we use constrained decoding to ensure that content outside of the stored documents is not generated. To increase speed, we only recall a short prefix in the second stage, and then locate its position to retrieve a complete passage. Experiments on KILT knowledge-sensitive tasks have verified that LLMs can independently recall reference passage location in various task forms, and the obtained reference significantly assists downstream tasks.
Generative Retrieval with Large Language Models
@@ 1 Introduction
Although LLMs have impressive capabilities, generating complete reference passage may be too time-consuming. To address this issue, we explore whether LLMs can locate passage by recalling a prefix only. First, we prompt the LLM to generate a short prefix, then use the FM-index to locate the document containing this prefix within the document set. Finally, we use the Knuth-Morris-Pratt (KMP) algorithm to identify the starting position of the prefix in the document, determining a longer passage as the final reference from this position. We find that this method significantly speeds up the LLM’s recall of reference. And the location of such a short prefix still maintains a high quality of reference retrieval. Our framework requires no additional retrieval models and pre-segmentation, allowing the LLM to independently and naturally obtain natural reference from articles of any length, ultimately naming it LLM2GR.
We conduct comprehensive experiments on 6 knowledge-sensitive tasks from the KILT benchmark (Petroni et al., 2021). We find that in a zero-shot setting, merely prompting the open-source LLMs Llama (Touvron et al., 2023a) and Llama-2 (Touvron et al., 2023b) enables the effective recall of needed page documents and fine-grained reference passage under the LLM2GR framework, enhancing the performance of downstream tasks. In summary, our main contributions are as follows:
@@
Figure 2: Shows the LLM2GR framework. First, all Wikipedia titles are stored in a prefix tree, then the LLM is prompted to recall title identifiers under this prefix tree constraint. Subsequently, an FM-index is constructed from the top $k$ documents obtained, and the LLM recalls reference passage under the new constraint. The grey area represents the short prefix, used for the next step of locating and extracting the complete passage.
2 LLM2GR
In this section, we detail our two-stage framework, LLM2GR. In the first stage, we prompt the LLM to recall title identifiers, which serve as candidate documents for the next stage. In the second stage, the LLM is prompted to recall reference passage from the documents obtained in the first stage. To increase speed, we only recall a short prefix, then locate and extract the reference within the document. The structure is shown in Figure 2.
@@ 2.3 Short Prefix Recall and Localization
Subsequently, we find the document $d$ corresponding to $p_{s}$ within the document set $D_{k}$ obtained in the first stage. Since we have controlled the number of documents in the document set $D_{k}$ obtained in the first stage, in most cases, we can obtain a unique document $d$ containing $p_{s}$. When we cannot obtain a unique document, we simply default to selecting the first document. In the next step, we use the Knuth-Morris-Pratt (KMP) string matching algorithm to quickly determine the starting position $st$ of $p_{s}$ in $d$, then extract a complete reference $p_{final}=d[st:st+l_{p}]$ starting from $st$, with a length of $l_{p}$ tokens. The autoregressive score of the short prefix is also calculated using Equation LABEL:eq2 and the final two-stage score is obtained using Equation LABEL:eq3 to select the final reference. In Experiment 3.2.1, we find that recalling just the prefix for localization still achieves good results in obtaining reference.
@@ 3.1 Experimental Setup
DPR 54.74 45.68 25.46 $\mathbf{16.19}$ 56.61 26.62
LLM2GR(Llama-7b) 54.46 $\mathbf{57.03}$ 44.56 15.13 76.57 52.91
LLM2GR(Llama-13b) 54.42 55.53 46.30 12.94 77.55 34.51
LLM2GR(Llama-2-7b) 56.33 56.43 46.20 14.60 77.29 49.61
LLM2GR(Llama-2-13b) $\mathbf{57.77}$ 54.41 $\mathbf{48.70}$ 15.00 $\mathbf{83.69}$ $\mathbf{57.63}$
@@
DPR $\mathbf{47.94}$ 66.60 20.29 14.40 41.22 45.38
LLM2GR(Llama-7b) 36.87 58.48 25.55 15.99 54.85 59.40
LLM2GR(Llama-13b) 37.72 60.96 26.34 14.80 55.20 50.79
LLM2GR(Llama-2-7b) 38.07 62.88 27.55 $\mathbf{16.85}$ 56.23 57.79
LLM2GR(Llama-2-13b) 40.82 $\mathbf{68.20}$ $\mathbf{30.04}$ 15.06 $\mathbf{58.42}$ $\mathbf{63.43}$
@@
DPR $\mathbf{33.49}$ 72.68 23.13 $\mathbf{20.75}$ 75.27 14.17
LLM2GR(Llama-7b) 29.78 70.18 24.61 20.60 78.10 14.47
LLM2GR(Llama-13b) 29.68 71.60 25.48 20.24 78.53 14.33
LLM2GR(Llama-2-7b) 29.89 70.04 25.55 20.50 78.04 14.48
LLM2GR(Llama-2-13b) 31.69 $\mathbf{72.94}$ 26.13 20.61 $\mathbf{78.79}$ $\mathbf{14.77}$
@@ 3.2.1 Page-level Results
Coarse-grained page-level results, as shown in Table 1, demonstrate that the LLM2GR framework, when implemented with Llama-2-13b, achieves the best R-precision scores of 57.77, 48.70, 83.69, and 57.63 on the NQ, HotpotQA, FEVER, and WoW datasets, respectively. This significantly surpasses the performance of sparse retrieval BM25 and dense retrieval Contriever in a zero-shot scenario. It also shows strong competitive power against the fully trained DPR method, especially on the WoW and FEVER datasets, with improvements of 27.08 and 31.01 points, respectively. This result is consistent with the hypothesis that LLMs are powerful in recalling coarse-grained title identifiers, enabling the acquisition of high-quality relevant pages that assist in the subsequent fine-grained recall stage.
@@ 3.2.2 Passage-level Results
Fine-grained reference passage results, as shown in Table 2, reveal that the LLM2GR framework, when implemented with Llama-2-13b, also achieves the best scores of 68.20, 30.04, 58.42, and 63.43 on the TriviaQA, HotpotQA, FEVER, and WoW datasets, respectively. We note that the improvement of the framework in fine-grained reference passage compared to the DPR method is relatively reduced compared to the page-level results. This suggests potential for optimization in activating LLMs to recall more detailed and longer reference, presenting a greater challenge compared to recalling shorter title. Notably, DPR performs excellently on the NQ dataset, which may relate to its training data format. Interestingly, in the HotpotQA dataset, BM25 remains competitive, surpassing dense retrieval methods, possibly due to the longer questions in this dataset leading to more vocabulary overlap. LLM2GR shows significant progress on the FEVER and WoW datasets, demonstrating the potential and adaptability of LLMs in recalling high-quality reference passage across different task formats. Furthermore, the general enhancement in performance with the progression from Llama to Llama-2 and the increase in model size indicates a correlation between the recall ability and the underlying capabilities of LLMs.
@@ 3.2.3 Downstream Task Results
Downstream task results are presented in Table 3. LLM2GR, based on Llama-2-13b recalled passage, achieved the best scores of 72.94, 78.79, and 14.77 on the TriviaQA, FEVER, and WoW downstream tasks, respectively, validating the performance of LLM recall references in downstream tasks. On the open-domain question answering NQ dataset, although DPR performed excellently after full data training, LLM2GR also displayed highly competitive performance. On the other hand, in the TriviaQA and HotpotQA datasets, due to the length of the questions, BM25 achieved excellent performance by obtaining more vocabulary overlap, yet LLM2GR still achieved comparable or better performance in most cases. The unsupervised trained Contriever performed relatively poorly across all tasks, emphasizing the crucial role of supervised training in enhancing the performance of dense retrieval models.
@@ 3.4 Further Analysis
Memory Usage Analysis Dense retrieval methods such as Contriever and DPR require over 60GB of memory usage. In contrast, sparse retrieval methods use far less memory, only needing 17GB. The LLM2GR framework, utilizing FM-index and Trie indexing, requires only 8GB when pre-encoding and storing all documents with FM-index, and the Trie storing all title identifiers needs just 25MB, which is negligible. Compared to sparse and dense retrieval methods, the recall framework effectively saves memory.
@@ 5 Conclusion and Future Work
This paper introduces a framework named LLM2GR, which utilizes LLMs to independently recall reference passage and can be flexibly applied to various knowledge-sensitive tasks. Mimicking the human habit of searching for information, we first prompt the LLM to recall relevant document pages, and then from these pages, recall and locate reference passage. Furthermore, through beam search constrained by Trie and FM-index, we ensure that the content recalled by the LLM is a subset of existing texts. This framework can be flexibly used with various open-source LLMs, expanding the potential applications of LLMs. In the future, we consider enhancing the ability of LLMs to recall under constraints through instruction tuning, as well as applying this framework to more retrieval fields, exploring lightweight methods to inject new document knowledge into LLMs, and integrating multi-hop reasoning into reference recall.
@@
Although LLM2GR demonstrates the potential of LLMs to recall reference passage in knowledge-sensitive tasks like humans, its application still faces several limitations. Firstly, this framework struggles to surpass the performance of the current state-of-the-art retrieval models, especially those models that have been fine-tuned on specific tasks through supervision. In the future, there is a need to explore more effective ways of instruction tuning for recalling under constraints. At the same time, LLM2GR relies on document title identifiers for phased recall, meaning that its recall capability may be limited for documents lacking clear titles or identifiers.
Moreover, the framework finds it challenging to effectively recall documents that appear less frequently in the pre-training stage. This indicates that if a document appears infrequently in the training data of the LLM, or if the document content significantly differs from the training data, LLM2GR may encounter difficulties in recalling these documents. For the updating of documents and the injection of new knowledge, LLM2GR requires additional training to inject this new information into the model parameters. There is still a need to explore more efficient, lightweight methods for injecting new documents in the future.
@@ C.1 Prompts for the First Stage
LLM2GR(Llama-7b) 54.46 57.03 44.56 36.87 58.48 25.55
LLM2GR(Vicuna-1.3-7b) 48.47 47.99 40.79 35.28 56.41 23.75
LLM2GR(Llama-13b) 54.42 55.53 46.30 37.72 60.96 26.34
LLM2GR(Vicuna-1.3-13b) 52.73 46.61 43.41 36.55 67.29 26.63
LLM2GR(Llama-2-7b) 56.33 56.43 46.20 38.07 62.88 27.55
LLM2GR(Llama-2-chat-7b) 3.31 1.12 0.98 4.09 3.97 3.43
LLM2GR(Vicuna-1.5-7b) 50.76 51.73 41.23 34.16 55.98 24.43
LLM2GR(Llama-2-13b) 57.77 54.41 48.70 40.82 68.20 30.04
LLM2GR(Llama-2-chat-13b) 1.94 1.60 1.55 6.38 7.93 4.71
LLM2GR(Vicuna-1.5-13b) 52.24 56.34 45.90 37.22 63.24 27.14
@@ Appendix D Experimental Results for Different Values of Alpha
In Figure 4, we compare the experimental results of LLM2GR when implemented based on Llama-2-13b with different $\alpha$ values. When $\alpha=0.0$, it’s equivalent to not having a two-stage weighted method, relying only on the scores from the second stage’s fine-grained passage recall, resulting in the selection of suboptimal reference. With the increase of $\alpha$, the model sees improvements in both page-level and passage-level results, proving the importance of the first stage document scores for the final reference selection. However, when $\alpha$ reaches 0.95 and continues to increase, the final performance actually decreases to some extent, indicating the need to find a balance between the two for better results.
@@ Appendix E Experimental Results for Different Prefix Lengths
In Figure 5, we conduct experiments with LLM2GR recalling different numbers of prefix tokens based on Llama-2-13b. We observe that longer prefix lengths do not bring additional performance improvements; on the contrary, they lead to a decrease in performance. Existing LLMs still perform better when generating shorter passages under constraints; longer passages introduce additional noise, resulting in decreased performance. However, overly short prefixes might also lack sufficient information, leading to an inability to accurately select the desired passage as a reference.
@@ Appendix G Experimental Results under Few-Shot Prompts
In Figure 6, we compare the experimental results of LLM2GR when incorporating few-shot prompts in the second stage of fine-grained passage recall based on Llama-2-13b. Using more sample prompts only brings partial improvements on the HotpotQA dataset, and the degree of improvement tends to stabilize when the number of samples increases from 3 to 5. On the NQ and TriviaQA datasets, no further improvements are observed, and a slight decline is even noted. There is still a need to explore more effective prompting methods.